OpenStudy (anonymous):

What is the simplified form of the expression: 3 log 5 + 3 log 2 – 2 log 4 ?

OpenStudy (anonymous):

3log5-log2

OpenStudy (anonymous):

Answer above is correct, if a little brief. In a little more detail\[3 \log 5 + 3 \log 2 – 2 \log 4=\log5^3+\log2^3-\log4^2=\log(\frac{125(8)}{16})=\log(\frac{125}{2})\]

OpenStudy (anonymous):

"Simplified form" might vary from teacher to teacher, and book to book, but we often consider a single log to be in simplest form.